OpenStudy (anonymous):

solve the systen 3x-2y=10 y=2x+5

OpenStudy (anonymous):

3x - 2y = 10 ________________ Add straight down 6x = 24 ...Divide by 6 x = 4 Substitute 4 in for x in one equation 3(4) + 2y = 14 12 + 2y = 14 ... Subtract 12 on both sides 2y = 2 y = 1

OpenStudy (angela210793):

3x-2(2x+5)=10 3x-4x=20 -x=20 x=-20 replace it to the 2nd eq y=-40+5 y=-35

OpenStudy (anonymous):

x=-20 y=-35